# Tensor Train Decomposition-based Channel Estimation for MIMO-AFDM Systems with Fractional Delay and Doppler

### 10 Mar 2026 v1

@@ Abstract
Affine Frequency Division Multiplexing (AFDM) has emerged as a promising chirp-based multicarrier technology for high-speed communication systems. By applying discrete affine Fourier transform (DAFT), AFDM has the potential to achieve significant improvements in spectral efficiency (SE) while maintaining a bit error rate (BER) performance comparable to that of OTFS. To fully exploit the diversity gain offered by AFDM, accurate channel estimation is essential. However, existing studies have mainly focused on the integer-delay-tap scenario and single-symbol pilot-based estimation. Since delay taps in practice are generally fractional, approximating them as integers not only degrades delay estimation accuracy but also severely affects Doppler frequency estimation. To address this problem, in this paper, we investigate channel estimation for multiple-input multiple-output (MIMO)-AFDM systems. A time-affine frequency (T-AF) domain pilot structure is proposed to exploit time-domain phase variations. By leveraging the rotational invariance property in the spatial and temporal domains, a channel estimation algorithm based on Vandermonde-structured tensor-train (TT) decomposition is developed. The proposed algorithm demonstrates superior computational efficiency compared with state-of-the-art parameter estimation methods. Moreover, diverging from current studies, we derive the global Ziv-Zakai bound (ZZB) as an alternative parameter estimation error lower bound to the Cramér-Rao bound (CRB). Numerical results show that the derived ZZB provides tighter global performance characterization and successfully captures the threshold phenomenon in mean square error (MSE) performance in the low-SNR regime. Furthermore, the proposed algorithm achieves superior communication performance relative to the existing schemes, while offering a computational speedup, reducing the execution time by an order of magnitude compared to the state-of-the-art iterative algorithms.
@@ I Introduction
To fundamentally overcome the performance degradation of OFDM in high-mobility scenarios, researchers have turned to novel modulation waveforms. One of the promising approaches is the delay-Doppler (DD) domain modulation, with orthogonal time frequency space (OTFS) being a representative technique [4, 5]. OTFS is a two-dimensional (2D) modulation scheme that transforms the multipath doubly selective channel into sparse and approximately time-invariant impulse responses in the delay-Doppler (DD) domain, thereby improving communication reliability. However, OTFS also suffers from notable drawbacks. Firstly, OTFS requires block-wise modulation over multiple symbols, which introduces inherent latency into the system [6]. Secondly, due to its 2D modulation structure, OTFS symbols exhibit intrinsic 2D coupling, preventing symbol-by-symbol processing as in OFDM. Instead, the receiver must perform joint detection over 2D signal blocks, which significantly increases detection complexity [7].
@@
  • Based on the proposed pilot structure, we model the received signal of MIMO-AFDM and formulate the channel estimation problem. By exploiting the rotational invariance property of MIMO-AFDM signals in space and time domains, as well as the impulse response magnitude characteristics in the affine frequency domain, we propose an efficient channel estimation algorithm based on Vandermonde-structured tensor-train (TT) decomposition. The proposed algorithm achieves significantly lower computational complexity than the existing iterative algorithms.
@@ III Pilot Structure Design And Channel Estimation Algorithm
In this section, we first introduce the proposed MIMO-AFDM embedded pilot structure and formulate the channel estimation problem. Then, by exploiting the multidimensional structure and inherent low-rank properties of the MIMO-AFDM signals, we develop a computationally efficient channel estimation framework based on Vandermonde-structured tensor-train (TT) decomposition tailored to the proposed pilot structure.
@@ III-A Proposed T-AF Embedded Pilot Structure
Let $m_{\mathrm{pilot}}$ and $N$ denote the position of the embedded pilot symbol in each AFDM symbols and the number of AFDM symbols that are inserted pilot symbols, respectively. By collecting all the pilot signals received on the $n_{\mathrm{BS}}$-th antenna, we have
@@ III-B1 Factor Matrices Retrieval
From (13) and (III-A), it can be seen that the tensor decomposition problem in (13) can be solved by canonical polyadic decomposition (CPD). Representative methods to solve such problem include a series of alternating iterative algorithms derived from the core ALS approach, such as the rectified ALS (RecALS) [20], the SVD-employed ALS [21] and tensorial minimum mean square error (T-MMSE) [22]. However, the robustness and convergence of ALS face significant challenges when the tensor rank exceeds 2 [23], and the computational efficiency of ALS based algorithms remains insufficient for high-mobility scenarios. As a novel tensor decomposition method, TT-decomposition [24] decomposes a high-dimensional tensor into multiple three-dimensional tensors connected in a chain structure, which offers greater flexibility and higher computational efficiency compared to CPD. Moreover, the decomposition error of the TT-decomposition can be strictly upper bounded by the sum of squared singular values discarded in the SVD of each dimension, effectively addressing the insufficient convergence and robustness of ALS. Without loss of generality, in the following derivations, we use $\bm{\mathcal{Y}}=\bm{\mathcal{S}}\times_{1}\mathbf{A}_{1}\times_{2}\mathbf{A}_{2}\times_{3}\mathbf{A}_{3}$ to denote the noiseless signal model and $\widehat{\bm{\mathcal{Y}}}$ to denote the noisy signal model.
The TT-decomposition represents tensor $\bm{\mathcal{Y}}$ as a chain-like product of three tensor cores as
@@
where $\mathbf{G}_{1}\in\mathbb{C}^{N_{\mathrm{BS}}\times P}$, $\bm{\mathcal{G}}_{2}\in\mathbb{C}^{P\times N\times P}$ and $\mathbf{G}_{3}\in\mathbb{C}^{P\times M_{\mathrm{region}}}$ denote the head matrix, the core tensor and the tail matrix, respectively. In this paper, the TT-SVD is applied to reconstruct the TT-cores $\mathbf{G}_{1}$, $\bm{\mathcal{G}}_{2}$ and $\mathbf{G}_{3}$ [25], which can be represented by
@@
where $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are the non-singular change-of-bias matrices. The singular value matrix is assumed to be incorporated into the right singular matrix. It is observed from the derivation that the CPP factor matrices are intrinsically embedded within the subspaces of the TT-cores. Specifically, the column spaces of $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ and $\mathbf{A}_{3}$ coincides with the column space of $\mathbf{G}_{1}$, the subspace spanned by the mode-2 fibers of $\bm{\mathcal{G}}_{2}$ and the row space of $\mathbf{G}_{3}$, respectively. Given the Vandermonde nature of $\mathbf{A}_{1}$ and $\mathbf{A}_{3}$, their defining parameters can be retrieved from the subspaces of $\mathbf{G}_{1}$ and the subspaces of the mode-2 unfolding of $\mathbf{G}_{2}$ via ESPRIT. Nevertheless, this approach incurs a column ambiguity. Since the parameters are estimated separately, there is no inherent mechanism to align the path indices, potentially resulting in mismatched pairs of channel parameters for the propagation paths.
@@ III-C Complexity Analysis
In this subsection, the computational of the proposed algorithm is analyzed. The main complexity of the proposed algorithm lies in SVD and pseudoinverse of matrices. Specifically, the TT-SVD process takes the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}\right)$. The EVD of (20) and the pseudoinverse and multiplication (III-B1) are on the order of the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{3}\right)$ and $\mathcal{O}\left(N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$, respectively. Finally, the total complexity of the proposed algorithm is of order $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}+N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$.
@@ V-B Communication Performance Comparison with Existing Methods
To evaluate the computational efficiency of different AFDM channel estimation algorithms, the average CPU running time versus SNR is demonstrated in Fig. 10. As shown in Fig. 10, the NOMP algorithm requires relatively high CPU time, as it needs to update the gradient matrix, the Hessian matrix, and the objective function at each iteration, all of which contribute to its high computational complexity. Furthermore, compared with NOMP, both the classical CP-ALS algorithm and the RecALS algorithm reduce the computational time by approximately one order of magnitude. This highlights the advantage of tensor decomposition-based methods in achieving low complexity when processing high-dimensional signals. Finally, since the proposed algorithm is based on Vandermonded structure TT-decomposition and involves no iterative process, the computational time of the proposed algorithm is further reduced by approximately one order of magnitude compared with CP-ALS and RecALS algorithm. The low computational complexity of the proposed algorithm makes it highly suitable for deployment in high-mobility scenarios.
@@
To further investigate the estimation performance of the algorithms under channels with different propagation scenarios, Fig. 11 shows the NMSE performance versus the number of propagation paths, where the SNR is set to 20 dB. It can be observed from Fig. 11 that the NMSE performance of the classical CP-ALS algorithm is mostly affected with the increase in the number of channel propagation paths. As the path number exceeds four, the NMSE performance of the CP-ALS deteriorates sharply. The main reason is the poor convergence behavior of thhe CP-ALS when processing the high-rank tensors [23]. Furthermore, the performance degradation of the RecALS algorithm is less pronounced as the number of paths increases. This is attributed to the fact that, unlike the classical CP-ALS, the RecALS fully leverages the Vandermonde factor matrices within the tensor. Finally, the proposed algorithm incurs less performance loss compared to the CP-ALS and the RecALS. In contrast to these two ALS-based algorithms, the proposed TT-decomposition-based algorithm demonstrates superior robustness in processing high-rank tensor signals.
@@ V-C Versus OTFS and OFDM
To further investigate the performance comparison between AFDM, OTFS and OFDM in more realistic environments, we simulate their BER and SE performance under conditions of fractional delay and Doppler. The maximum velocity is set to 500 km/h, the total number of subcarriers is set to $M=512$, and the number of symbols in the OTFS frame is $N=13$, resulting in a maximum Doppler tap of $3.43$. According to [36], the guard interval is equal to the ceiling of the maximum Doppler tap. This implies that the pilot length in the time domain extends to $N$ (i.e., the full frame duration). The number of pilots used for AFDM is set to $N_{p}=7$  [^1] For the proposed algorithm, it is only necessary when the number of pilot symbols satisfies $N_{p}\geq P$.. The other parameters are the same as in Table I.
@@ VI Conclusion
In this paper, we investigated the channel estimation for MIMO-AFDM systems and analyzed the input-output relationship for MIMO-AFDM systems with the consideration of fractional delay and Doppler. To tackle with the estimation of fractional delay and Doppler, we proposed a T-AF domain embedded pilot structure. By leveraging the magnitude of AF domain and the rotational invariance property of time domain, we proposed an efficient Vandermonde structure TT decomposition-based channel estimation algorithm. Compared to the existing parameter algorithms, the proposed algorithm significantly reduce the computational complexity. To evaluate the MSE performance, we derived Ziv-Zakai bound for the 3D MIMO-AFDM signal model. Simulation results showed that the proposed algorithm significantly outperforms the benchmarks at medium-to-high SNR.

### 20 Mar 2026 v2

@@ Abstract
Affine Frequency Division Multiplexing (AFDM) has emerged as a promising chirp-based multicarrier technology for high-speed communication systems. By applying discrete affine Fourier transform (DAFT), AFDM has the potential to achieve significant improvements in spectral efficiency (SE) while maintaining a bit error rate (BER) performance comparable to that of OTFS. To fully exploit the diversity gain offered by AFDM, accurate channel estimation is essential. However, existing studies have mainly focused on the integer-delay-tap scenario and single-symbol pilot-based estimation. Since delay taps in practice are generally fractional, approximating them as integers not only degrades delay estimation accuracy but also severely affects Doppler frequency estimation. To address this problem, in this paper, we investigate channel estimation for multiple-input multiple-output (MIMO)-AFDM systems. A time-affine frequency (T-AF) domain pilot structure is proposed to exploit time-domain phase variations. By leveraging the rotational invariance property in the spatial and temporal domains, a channel estimation algorithm based on Vandermonde-structured tensor-train ($T$) decomposition is developed. The proposed algorithm demonstrates superior computational efficiency compared with state-of-the-art parameter estimation methods. Moreover, diverging from current studies, we derive the global Ziv-Zakai bound (ZZB) as an alternative parameter estimation error lower bound to the Cramér-Rao bound (CRB). Numerical results show that the derived ZZB provides tighter global performance characterization and successfully captures the threshold phenomenon in mean square error (MSE) performance in the low-SNR regime. Furthermore, the proposed algorithm achieves superior communication performance relative to the existing schemes, while offering a computational speedup, reducing the execution time by an order of magnitude compared to the state-of-the-art iterative algorithms.
@@ I Introduction
To fundamentally overcome the performance degradation of OFDM in high-mobility scenarios, researchers have turned to novel modulation waveforms. One of the promising approaches is the delay-Doppler ($D$) domain modulation, with orthogonal time frequency space (OTFS) being a representative technique [4, 5]. OTFS is a two-dimensional (2D) modulation scheme that transforms the multipath doubly selective channel into sparse and approximately time-invariant impulse responses in the delay-Doppler ($D$) domain, thereby improving communication reliability. However, OTFS also suffers from notable drawbacks. Firstly, OTFS requires block-wise modulation over multiple symbols, which introduces inherent latency into the system [6]. Secondly, due to its 2D modulation structure, OTFS symbols exhibit intrinsic 2D coupling, preventing symbol-by-symbol processing as in OFDM. Instead, the receiver must perform joint detection over 2D signal blocks, which significantly increases detection complexity [7].
@@
  • Based on the proposed pilot structure, we model the received signal of MIMO-AFDM and formulate the channel estimation problem. By exploiting the rotational invariance property of MIMO-AFDM signals in space and time domains, as well as the impulse response magnitude characteristics in the affine frequency domain, we propose an efficient channel estimation algorithm based on Vandermonde-structured tensor-train ($T$) decomposition. The proposed algorithm achieves significantly lower computational complexity than the existing iterative algorithms.
@@ III Pilot Structure Design And Channel Estimation Algorithm
In this section, we first introduce the proposed MIMO-AFDM embedded pilot structure and formulate the channel estimation problem. Then, by exploiting the multidimensional structure and inherent low-rank properties of the MIMO-AFDM signals, we develop a computationally efficient channel estimation framework based on Vandermonde-structured tensor-train ($T$) decomposition tailored to the proposed pilot structure.
@@ III-A Proposed T-AF Embedded Pilot Structure
Let $m_{\mathrm{pilot}}$ and NN denote the position of the embedded pilot symbol in each AFDM symbols and the number of AFDM symbols that are inserted pilot symbols, respectively. By collecting all the pilot signals received on the $n_{\mathrm{BS}}$-th antenna, we have
@@ III-B1 Factor Matrices Retrieval
From (13) and (III-A), it can be seen that the tensor decomposition problem in (13) can be solved by canonical polyadic decomposition (CPD). Representative methods to solve such problem include a series of alternating iterative algorithms derived from the core ALS approach, such as the rectified ALS (RecALS) [20], the SVD-employed ALS [21] and tensorial minimum mean square error (T-MMSE) [22]. However, the robustness and convergence of ALS face significant challenges when the tensor rank exceeds 2 [23], and the computational efficiency of ALS based algorithms remains insufficient for high-mobility scenarios. As a novel tensor decomposition method, $T$-decomposition [24] decomposes a high-dimensional tensor into multiple three-dimensional tensors connected in a chain structure, which offers greater flexibility and higher computational efficiency compared to CPD. Moreover, the decomposition error of the $T$-decomposition can be strictly upper bounded by the sum of squared singular values discarded in the SVD of each dimension, effectively addressing the insufficient convergence and robustness of ALS. Without loss of generality, in the following derivations, we use $\bm{\mathcal{Y}}=\bm{\mathcal{S}}\times_{1}\mathbf{A}_{1}\times_{2}\mathbf{A}_{2}\times_{3}\mathbf{A}_{3}$ to denote the noiseless signal model and $\widehat{\bm{\mathcal{Y}}}$ to denote the noisy signal model.
The $T$-decomposition represents tensor $\bm{\mathcal{Y}}$ as a chain-like product of three tensor cores as
@@
where $\mathbf{G}_{1}\in\mathbb{C}^{N_{\mathrm{BS}}\times P}$, $\bm{\mathcal{G}}_{2}\in\mathbb{C}^{P\times N\times P}$ and $\mathbf{G}_{3}\in\mathbb{C}^{P\times M_{\mathrm{region}}}$ denote the head matrix, the core tensor and the tail matrix, respectively. In this paper, the $T$-SVD is applied to reconstruct the $T$-cores $\mathbf{G}_{1}$, $\bm{\mathcal{G}}_{2}$ and $\mathbf{G}_{3}$ [25], which can be represented by
@@
where $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are the non-singular change-of-bias matrices. The singular value matrix is assumed to be incorporated into the right singular matrix. It is observed from the derivation that the CPP factor matrices are intrinsically embedded within the subspaces of the $T$-cores. Specifically, the column spaces of $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ and $\mathbf{A}_{3}$ coincides with the column space of $\mathbf{G}_{1}$, the subspace spanned by the mode-2 fibers of $\bm{\mathcal{G}}_{2}$ and the row space of $\mathbf{G}_{3}$, respectively. Given the Vandermonde nature of $\mathbf{A}_{1}$ and $\mathbf{A}_{3}$, their defining parameters can be retrieved from the subspaces of $\mathbf{G}_{1}$ and the subspaces of the mode-2 unfolding of $\mathbf{G}_{2}$ via ESPRIT. Nevertheless, this approach incurs a column ambiguity. Since the parameters are estimated separately, there is no inherent mechanism to align the path indices, potentially resulting in mismatched pairs of channel parameters for the propagation paths.
@@ III-C Complexity Analysis
In this subsection, the computational of the proposed algorithm is analyzed. The main complexity of the proposed algorithm lies in SVD and pseudoinverse of matrices. Specifically, the $T$-SVD process takes the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}\right)$. The EVD of (20) and the pseudoinverse and multiplication (III-B1) are on the order of the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{3}\right)$ and $\mathcal{O}\left(N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$, respectively. Finally, the total complexity of the proposed algorithm is of order $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}+N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$.
@@ V-B Communication Performance Comparison with Existing Methods
To evaluate the computational efficiency of different AFDM channel estimation algorithms, the average CPU running time versus SNR is demonstrated in Fig. 10. As shown in Fig. 10, the NOMP algorithm requires relatively high CPU time, as it needs to update the gradient matrix, the Hessian matrix, and the objective function at each iteration, all of which contribute to its high computational complexity. Furthermore, compared with NOMP, both the classical CP-ALS algorithm and the RecALS algorithm reduce the computational time by approximately one order of magnitude. This highlights the advantage of tensor decomposition-based methods in achieving low complexity when processing high-dimensional signals. Finally, since the proposed algorithm is based on Vandermonded structure $T$-decomposition and involves no iterative process, the computational time of the proposed algorithm is further reduced by approximately one order of magnitude compared with CP-ALS and RecALS algorithm. The low computational complexity of the proposed algorithm makes it highly suitable for deployment in high-mobility scenarios.
@@
To further investigate the estimation performance of the algorithms under channels with different propagation scenarios, Fig. 11 shows the NMSE performance versus the number of propagation paths, where the SNR is set to 20 dB. It can be observed from Fig. 11 that the NMSE performance of the classical CP-ALS algorithm is mostly affected with the increase in the number of channel propagation paths. As the path number exceeds four, the NMSE performance of the CP-ALS deteriorates sharply. The main reason is the poor convergence behavior of thhe CP-ALS when processing the high-rank tensors [23]. Furthermore, the performance degradation of the RecALS algorithm is less pronounced as the number of paths increases. This is attributed to the fact that, unlike the classical CP-ALS, the RecALS fully leverages the Vandermonde factor matrices within the tensor. Finally, the proposed algorithm incurs less performance loss compared to the CP-ALS and the RecALS. In contrast to these two ALS-based algorithms, the proposed $T$-decomposition-based algorithm demonstrates superior robustness in processing high-rank tensor signals.
@@ V-C Versus OTFS and OFDM
To further investigate the performance comparison between AFDM, OTFS and OFDM in more realistic environments, we simulate their BER and SE performance under conditions of fractional delay and Doppler. The maximum velocity is set to 500 km/h, the total number of subcarriers is set to $M=512$, and the number of symbols in the OTFS frame is $N=13$, resulting in a maximum Doppler tap of $3.43$. According to [36], the guard interval is equal to the ceiling of the maximum Doppler tap. This implies that the pilot length in the time domain extends to NN (i.e., the full frame duration). The number of pilots used for AFDM is set to $N_{p}=7$  [^1] For the proposed algorithm, it is only necessary when the number of pilot symbols satisfies $N_{p}\geq P$.. The other parameters are the same as in Table I.
@@ VI Conclusion
In this paper, we investigated the channel estimation for MIMO-AFDM systems and analyzed the input-output relationship for MIMO-AFDM systems with the consideration of fractional delay and Doppler. To tackle with the estimation of fractional delay and Doppler, we proposed a T-AF domain embedded pilot structure. By leveraging the magnitude of AF domain and the rotational invariance property of time domain, we proposed an efficient Vandermonde structure $T$ decomposition-based channel estimation algorithm. Compared to the existing parameter algorithms, the proposed algorithm significantly reduce the computational complexity. To evaluate the MSE performance, we derived Ziv-Zakai bound for the 3D MIMO-AFDM signal model. Simulation results showed that the proposed algorithm significantly outperforms the benchmarks at medium-to-high SNR.